### v1

# August 11, 2025

Abstract
In this short-length paper, we will present some math that play a central role in quantum hydrodynamics and were presented by Mita in the past. In this formulation of QM, a quantity is involved which is the diffusion flux. As we will see, this quantity plays a crucial role on the evolution of wave packets. More specifically, we will initially briefly explain some results derived in Mita’s papers, by analyzing Gaussian and soliton wave packets, while our main focus is on the role of diffusion in the evolution of super-luminal wave packets.
I Introduction
In mita2003dispersive Mita had introduced the quantum momentum density which contains a probability flux $\overrightarrow{J}$ term and a diffusion term $\overrightarrow{D}$
$\Psi^{*}(r,t)\hat{P}\Psi(r,t)=M\left(\overrightarrow{J}(r,t)+i\overrightarrow{D}(r,t)\right)$
where $M$ is the mass of the quantum particle. Now, using $\hat{P}=-i\hbar\overrightarrow{\nabla}$, where $\hbar$ is the Planck’s constant, we obtain:
$\overrightarrow{J}(r,t)=\frac{\hbar}{M}Im\left(\Psi^{*}(r,t)\overrightarrow{\nabla}\Psi(r,t)\right)$ (1)
and
$\overrightarrow{D}(r,t)=\frac{\hbar}{M}Re\left(\Psi^{*}(r,t)\overrightarrow{\nabla}\Psi(r,t)\right)$
Using $\Psi(r,t)=R(r,t)e^{iS(r,t)}$, we are able to derive:
$\overrightarrow{D}(r,t)=-\frac{\hbar}{2M}\overrightarrow{\nabla}\rho(r,t)$ (2)
which is also the expression for the diffusion flux in classical diffusion. Mita also showed that both the probability and diffusion flux contribute to the total kinetic energy of the particle:
$E_{K}=\frac{1}{2}M\left(\int_{V}{\frac{J^{2}(r,t)}{\rho(r,t)}}d\tau+\int_{V}{\frac{D^{2}(r,t)}{\rho(r,t)}}d\tau\right)$
In contrast to Bohm’s theory bohm1952suggested , where the particle doesn’t move in stationary states (suppose a 1D system) since $\overrightarrow{v}^{\Psi}=\frac{\overrightarrow{J}^{\Psi}}{\rho}=0$, here we have quantum fluctuations that lead to a quantum Brownian motion. In other words, there is still contribution to the total kinetic energy due to diffusion. This kinetic energy is associated with the momentum uncertainty of the particle due to fluctuations. More specifically:
$E_{K,x}^{(stationary)}=\frac{(\Delta P)^{2}}{2M}=\frac{1}{2}M\int_{\mathbb{R}}{\frac{D^{2}(x)}{\rho(x)}}dx$
The diffusion flux is an indicator of the tendency of the wavepacket to diffuse. We should note also that, in the past, there was also a quantity named osmotic velocity that was discussed by Bohm and Hiley bohm1989non in the context of a stochastic version of Bohmian mechanics. This velocity was defined as follows:
$\overrightarrow{u}(r,t)=\frac{\hbar}{2M}\frac{\overrightarrow{\nabla}\rho(r,t)}{\rho(r,t)}=-\frac{\overrightarrow{D}(r,t)}{\rho(r,t)}$
The role of that velocity according to the authors was to balance diffusion in stationary state systems, where there was no probability evolution over time. This velocity, according to the authors, is responsible for ”guiding” the particle in higher probability areas. In addition, it was noted that the osmotic velocity becomes infinite at the nodes of the wave function. We have to add that we know that for every position square-integrable wave function $\Psi(x)$, we have a corresponding momentum wave function $\Psi(\rho)$ which is also square-integrable. This means that $|\Psi(\rho\rightarrow\infty)|^{2}=0$. In other words, it’s very difficult to find a particle that has extremely large velocity. And indeed, this is what we observe with the osmotic velocity. The particle has very large velocity close to the nodes of the wave function, where it is obviously very difficult to be observed. And that, of course, sounds very natural since the dwell time of a particle at a specific spatial region is inversely proportional to its velocity in that region.
Now that we are finished with the osmotic velocity, we will move on to the analysis of the evolution of wave packets. In chapter II we will briefly discuss both Gaussian and soliton wave packets and at the end, in Chapter III, we will see the role of diffusion flux in the evolution of initially extremely localized quantum systems.
II DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
The role of this Chapter is not to produce new knowledge, but to introduce the diffusion flux by looking specifically at the evolution of two types of waveforms. First, we have the Gaussian wave packet:
$\Psi_{g}(x,t)=\left(\frac{2}{\pi}\right)^{\frac{1}{4}}\frac{1}{\sqrt{\pi}}e^{-(x-u_{0}t)^{2}/\epsilon^{2}}e^{i(k_{0}x-\frac{\hbar k_{0}^{2}}{2m}t+\delta)}$
where $\epsilon(t)=\alpha\sqrt{1+\left(\frac{t}{T}\right)^{2}}$ , $u_{0}=\frac{\hbar k_{0}}{m}$ , $T=\frac{ma^{2}}{2\hbar}$ and
$\delta=\frac{(x-u_{0}t)^{2}}{\epsilon^{2}}\frac{t}{T}-\frac{1}{2}tan^{-1}\left(\frac{t}{T}\right)$
If we use Eqs. (1) and (2) according to Mita mita2021schrodinger we get
$\overrightarrow{J}_{g}(x,t)=\rho_{g}(x,t)u_{0}\overrightarrow{e}_{x}+\frac{mt}{T}\overrightarrow{D}_{g}(x,t)$
The first term is related to the ”centre of mass” motion of the wave packet and the second term, which contains the diffusion flux is related to the expansion of the probability distribution. Now, when it comes to the soliton wave packet, Mita mita2021schrodinger defined a probability distribution:
$\rho_{sol}(x,t)=\frac{1}{2\sigma_{0}}sech^{2}\left(\frac{x-u_{0}t}{\sigma_{0}}\right)$
and found the potential $V(x,t)$ so that $\rho_{sol}$ behaves as a soliton. He found that:
$V(x,t)=-\frac{2\hbar^{2}}{m\sigma_{0}}\rho_{sol}(x,t)$
. At first sight, from a hydrodynamical perspective, this equation doesn’t give us any physical intuition. However, this equation can be written in the following form:
$\overrightarrow{F}(x,t)=\frac{2\hbar^{2}}{m\sigma_{0}}\overrightarrow{\nabla}\rho_{sol}(x,t)\sim-\overrightarrow{D}_{sol}(x,t)$
And that’s how the form of the wave packet is preserved. The greater the tendency of the wave packet to diffuse (greater $\overrightarrow{D}$) the larger the external force of opposite direction that has to be applied. And, of course, the coefficient $\mu=\frac{2\hbar^{2}}{m\sigma_{0}}$ is position independent, because otherwise it would cause deformation of the wave packet.
III DIFFUSION AND SUPER-LUMINAL WAVE PACKETS
In this Chapter, we will examine a state that at $t=0$ is extremely localized. More specifically, we wiil consider the initial state:
$\Psi(x,0)=\frac{1}{\sqrt{a}}\chi_{[-a/2,a/2]}(x)e^{ik_{0}x}$
According to Mita again mita2007dispersion the time-evolved state in free space will be the following:
$\Psi(x,t)=\frac{(-1)^{3/4}}{2\sqrt{ia}}\left[erfi\left[(-1)^{1/4}\sqrt{\frac{m}{2\pi\hbar t}}(x-a/2)\right]-erfi\left[(-1)^{1/4}\sqrt{\frac{m}{2\pi\hbar t}}(x+a/2)\right]\right]$
After very short time and very far very away from $x=a/2$, we have that:
$\rho(x>>a/2,t<<m/2\pi\hbar)\sim\left|\frac{e^{z_{1}^{2}}}{\sqrt{\pi}z_{1}}\left(1+\frac{1}{2z_{1}^{2}}+O\left(\frac{1}{z_{1}^{4}}\right)\right)-\frac{e^{z_{2}^{2}}}{\sqrt{\pi}z_{2}}\left(1+\frac{1}{2z_{2}^{2}}+O\left(\frac{1}{z_{2}^{4}}\right)\right)\right|^{2}$
where $z_{1}=\sqrt{\frac{m}{2i\hbar t}}(x-a/2)$ and $z_{2}=\sqrt{\frac{m}{2i\hbar t}}(x+a/2)$ From above, we can conclude that
$Prob(x\in[L_{1},L_{2}],t<<m/2\pi\hbar|L_{2},L_{1}>>a/2)\approx\int_{L_{1}}^{L_{2}}\rho(x>>a/2,t<<m/2\pi\hbar)dx\neq 0$
This reveals a non-zero probability of finding the particle infinitely far away from $x=a/2$ after very small time. But, hydrodynamically speaking, what is the cause of that infinite diffusion of the wave packet? The answer is the infinite diffusion flux for the initial state at the points x=a/2x=-a/2 and $x=a/2$, that reveals the infinite diffusion tendency of the initial wave packet at those points. The infinite suppression of the initial wave packet causes enormous diffusive tendency and this manifests itself after the removal of the physical ”walls” that restrict the wave packet. More specifically, we have:
$\Psi(x,0)=\frac{1}{\sqrt{a}}\Theta(a/2-|x|)e^{ik_{0}x}=\frac{1}{\sqrt{a}}\Theta(a/2-x)\Theta(a/2+x)e^{ik_{0}x}$
By applying Eq. (2) to the system and given that $\frac{d\Theta(x)}{dx}=\delta(x)$, as well as $\Theta^{2}(x)=\Theta(x)$ we obtain the expression:
$\overrightarrow{D}(x,0)=-\frac{\hbar}{2ma}\left[\delta(a/2+x)\Theta(a/2-x)-\delta(a/2-x)\Theta(a/2+x)\right]\overrightarrow{e}_{x}$
where $\overrightarrow{e}_{x}$ a unit vector that points to the positive $x$-direction. We observe that
$\lim_{x\rightarrow a/2}\overrightarrow{D}(x,0)=(+\infty)\overrightarrow{e}_{x}$
and
$\lim_{x\rightarrow-a/2}\overrightarrow{D}(x,0)=(-\infty)\overrightarrow{e}_{x}$
as expected.
IV DISCUSSIONS AND CONCLUSIONS
In this paper, we managed to give some clues to support the idea of contribution of the osmotic velocity field on quantum dynamics, even on stationary states. We also presented the mathematical and physical/hydrodynamical reason behind the infinite diffusion of strictly localized wave packets. In conclusion, if someone wants to examine the evolution of a wave-packet they’d better look at the Bohmian velocity. However, as we saw, diffusion flux also provides us with significant information relevant to the origin of quantum diffusion, which arises from the diffusion tendencies of the initial state (that are expressed by the diffusion flux).